OpenStudy (anonymous):

Evaluate: ln2 + 2ln3 - ln18

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

ln2 + 2ln3 - ln18 =ln[2*3^2/18] =ln1 = 0 lnx= ln4+ln7 lnx=ln(4*7) lnx=ln28 x=28

OpenStudy (anonymous):

sorry it is your 1st answer omits second oart

OpenStudy (anonymous):

Thank you @irshadnol :) Helped me alot :)